# HGT: Leveraging Heterogeneous Graph-enhanced Large Language Models for Few-shot Complex Table Understanding

### v1

@@ Abstract
Table understanding (TU) has achieved promising advancements, but it faces the challenges of the scarcity of manually labeled tables and the presence of complex table structures. To address these challenges, we propose HGT, a framework with a heterogeneous graph (HG)-enhanced large language model (LLM) to tackle few-shot TU tasks. It leverages the LLM by aligning the table semantics with the LLM’s parametric knowledge through soft prompts and instruction turning and deals with complex tables by a multi-task pre-training scheme involving three novel multi-granularity self-supervised HG pre-training objectives. We empirically demonstrate the effectiveness of HGT, showing that it outperforms the SOTA for few-shot complex TU on several benchmarks.
HGT: Leveraging Heterogeneous Graph-enhanced Large Language Models for Few-shot Complex Table Understanding
Rihui Jin1, Yu Li1, Guilin Qi1, Nan Hu1, Yuan-Fang Li2, Jiaoyan Chen3, Jianan Wang4, Yongrui Chen1, Dehai Min1 1Southeast University 2Monash University 3The University of Manchester 4Alibaba Group {jrh, yuli_11, gqi, nanhu, yrchen, zhishanq}@seu.edu.cn YuanFang.Li@monash.edujiaoyan.chen@manchester.ac.uk nanxi.wjn@alibaba-inc.com
@@ 1 Introduction
Table Understanding (TU) seeks to learn informative embeddings of tables containing inherently tabular semantics, often formatted in ways not easily interpretable by machines as shown in Fig. 1. This endeavor enhances machine performance across a range of table-related tasks, such as Table QA Herzig et al. (2020) and Cell Type Classification Ghasemi-Gol et al. (2019) and Table Type Classification Wang et al. (2020b).
@@
Yet, in real-world scenarios, TU faces the challenges of a lack of sufficient human annotations and the presence of complex table structures, which diminishes the effectiveness and applicability of existing frameworks. As for the first challenge, the data-hungry nature of existing frameworks results in diminished performance in few-shot TU scenarios, where only several samples are annotated. Although some studies Yin et al. (2020); Herzig et al. (2020) have adopted pre-training with encoder-only architectures Devlin et al. (2019) to alleviate the annotation burden, these solutions still require considerable amounts of labeled data for task-specific fine-tuning. As for the second challenge, while current frameworks attempt to capture structural information in tables through the use of position embeddings Wang et al. (2020b) or by modeling tables as graphs Wang et al. (2021), these methods are effective for simple tables but wane with complex tables. This shortfall arises because the cell-to-cell relationships in complex tables are more intricate than those encountered in simple tables as shown in Fig. 1.
@@
Figure 1: Few-shot complex Table Understanding. Complex tables contain intricate cell-to-cell relationships including dependency, hierarchical, and parallel ones.
@@
Fortunately, recent advancements have introduced techniques that can be integrated into the TU framework to tackle the challenges: 1) Multimodal Large Language Models (LLMs) have shown remarkable effectiveness in managing few-shot tasks involving data from other modalities, such as vision Liu et al. (2023) and time series Jin et al. (2023a). This is achieved through instruction tuning and soft prompts, which seek to align the semantic spaces of the other modality encoder with the LLM. 2) Self-supervised heterogeneous graph (HG) pre-training Ma et al. (2023); Yang et al. (2022) empowers models to navigate multifaceted relationships within data from substantial volumes of unlabeled data.
@@
In this work, we propose HGT, an HG-enhanced LLM framework for the few-shot complex TU. To enable HGT to model the intricate relationships within complex tables, HGT begins with modeling the tabular data with the HG and then processing it through a tabular graph encoder to generate vectors imbued with structural information. Additionally, to ensure the LLM achieves comparable performance in table tasks to its performance in natural language (NL) tasks, especially in few-shot scenarios, we align the representation spaces of the tabular graph encoder and the LLM via instruction tuning. Specifically, we incorporate table vectors containing structural information as soft prompts within the LLM inputs and innovatively design three multi-granularity self-supervised tasks tailored for both tables and LLMs for pre-training. After pre-training specific parameters of HGT via these self-supervised tasks, the HGT capitalizes on the LLM’s exceptional generalization capability to adapt to downstream tasks with minimal data samples. To validate the performance of HGT in few-shot complex TU, we conduct extensive comparative experiments with existing powerful baselines on eight publicly available datasets for three specific table-related tasks in few-shot scenarios. The experimental results show that HGT exceeds the current SOTA for few-shot complex TU across multiple benchmark datasets.
In summary, our main contributions are:
@@
  • To the best of our knowledge, we are the first to propose a framework to align table modality with NLs to empower the LLM to perform on tables with the same few-shot proficiency as it does with NLs.
@@
  • We conduct extensive experiments on eight publicly available datasets, and the experimental results show that HGT exceeds the current SOTA for few-shot complex TU across multiple benchmark datasets.
@@ 2.1 Graph-based Table Understanding
Many studies Du et al. (2021); Jin et al. (2023b) have tried to convert tables into graphs and utilized graph encoders to capture tables’ inherent topological information. These frameworks include employing homogeneous graphs or utilizing a basic node-linking strategy that connects cells exclusively to their adjacent counterparts. Consequently, such frameworks underperform when dealing with complex table structures.
2.2 LLM-based Table Understanding
Following the remarkable success of LLMs in NL tasks, some efforts Zhang et al. (2023); Chen (2022) have extended their application to table-related tasks. Although these frameworks leverage the exceptional generalization capacity of LLMs to achieve SOTA performance in some few-shot tasks, they resort to simply converting the table into a row-by-row NL format for input into LLMs. This process leads to a loss of the tables’ intrinsic topological information.
2.3 Query Statement-based Table QA
Several studies Cao et al. (2023); Cheng et al. (2022) have adapted semantic parsing techniques, traditionally applied to database tables, to general Table QA tasks. This involves transforming the table into a format interpretable by query languages, such as SQL, and subsequently utilizing Codex to generate a query statement to retrieve the answer. This methodology represents the current SOTA in Table QA. However, its architectural limitations restrict its applicability to other table-related tasks.
@@ 2.4 Bert-like Encoder-only Table Understanding
Since the rise in popularity of pre-training models like BERT Devlin et al. (2019), there has been considerable effort devoted to designing specialized encoding methods for tabular data and unique pre-text objectives for pre-training Herzig et al. (2020); Cheng et al. (2021a); Jia et al. (2023). Despite their utilization of self-supervised training, these methods still require a substantial amount of labeled data during fine-tuning for downstream tasks. Additionally, while they incorporate positional embedding into serialized tabular data, they do not effectively capture topological information.
@@ 3 Task Definition
Given a table $T=\{c_{i,j}|0\leq i<N,0\leq j<M\}$ where $N$ is the number of rows, $M$ is the number of columns, and $c_{i,j}$ is the cell located in the $i^{th}$ row and $j^{th}$ column. Merged cells, characterized by a row span or column span greater than 1, are prevalent in complex tables. For instance, the cell labeled "project" in the top-left corner of the table shown in Fig. 3 has a row span of 2. We assign the coordinates of such merged cells based on the position of the top-left cell before merging. Hence, the coordinate of the "project" cell is designated as (0,0).
Sub-TU tasks for Evaluation
@@
Table QA demands that the model produce an answer $y_{a}$ in response to a natural language question $q$, with table $T$ serving as the reference for deriving the answer $y_{a}$.
@@ 4 METHODOLOGY
Figure 2: An overview of HGT framework. HGT processes <table, instruction> as an input. First, the table is converted into an HG and processed by a Tabular HG encoder to generate vectors for each tabular node, while the LLM transforms instruction texts into initial token embeddings. Subsequently, the HG encoder’s outputs serve as soft prompts for the LLM, enabling the replacement of placeholder embeddings with actual tabular node vectors. The modified embedding sequence is then processed by the remaining LLM layers. Throughout both Stage 1 and Stage 2, only the weights of red components are tuned.
This section explains the process in three phases: tabular HG construction and two tuning stages. An overview of HGT is shown in Fig. 2.
@@ 4.1 Tabular Heterogeneous Graph Construction
Figure 3: Converting a table into a heterogeneous graph. Cell nodes of varying types are depicted using distinct colors: Table nodes in green, Row nodes in red, Data Cell nodes in blue, and Header Cell nodes in yellow. Similarly, edges are differentiated by color to represent various types. Lines without arrows indicate bidirectional edges. For the sake of clarity, some edges have been omitted.
@@
Given that HGs are more proficient at capturing diverse relationships compared to homogeneous graphs, we employ HGs in conjunction with a refined method that considers the semantic roles of cells to effectively model the structure of complex tables. The subsequent subsections detail the process of converting tabular data into HGs, including the creation of nodes and the heuristic rules for establishing edges between nodes. The process of creating a tabular HG is shown in Fig. 3.
4.1.1 Tabular Node Construction
@@
Four Tabular Node Types. TABLE, ROW, Header CELL, and Data CELL nodes are denoted as green, red, yellow, and blue nodes in Fig. 3. 1) The TABLE node represents the content described by the table, facilitating the table-level tasks. 2) The ROW node signifies the information contained within a row, aiding in the prediction of the row’s type during self-supervised training. 3) The Header CELL node denotes cells located in the header row, identifying column schemas or categories. 4) The Data CELL node represents cells in data rows, meaning the actual data entries of the table.
@@
Creating Tabular Nodes. First, CELL nodes are created for each cell in the table, with each node denoted as $c_{i,j}$, where $i$ and $j$ represent the cell’s coordinates in the original table. If a CELL node is located in the table’s Top Header Row or is part of a merged cell spanning the entire table width, it is viewed as a Header Cell; otherwise, it is identified as a Data Cell. The method for determining the Top Header Row Number of a table is elaborated in Appendix B. Subsequently, an equivalent number of ROW nodes is created to correspond with the number of rows in the table, along with one TABLE node.
@@
Initializing Tabular Node Embeddings. An initialization vector is required for each node in the HG. For Header and Data CELL nodes, we employ the output of Sentence-BERT Reimers and Gurevych (2019) applied to the text within each cell to obtain their initialization vectors. In the case of ROW nodes $r_{i}$, we initialize a vector by concatenating the text from the cells in the $i^{th}$ row and inputting this concatenated text into the Sentence-BERT. TABLE nodes serve to represent the table’s content itself. In a human’s view, a table’s content can be inferred by examining the cells in the headers. Consequently, we opt to concatenate the text from the cells in the headers and to obtain the embedding.
4.1.2 Adding Edges
@@
To enhance the machine’s understanding of the table’s semantics, we analyze complex table data and develop the following heuristic method to link nodes: 1) A Table node, representing the whole table, should be linked to all CELL nodes to encapsulate the global semantics. 2) A Row node derives information from cells within the same row, so the Row node $r_{i}$ should be linked to each Cell node $c_{i*}$. 3) A strong correlation exists between the semantics of data cells and their corresponding header cells within the same column. 4) Data cells located in the same column exhibit a stronger relational bond compared to those in different columns. Consequently, adjacent data cells within a column are interconnected. 5) The interpretation of relationships between data cells across columns necessitates the semantic understanding of their respective header cells. For instance, as demonstrated in Fig. 3, determining the relationship between cell $c_{2,0}$ and cell $c_{2,1}$ requires an examination of their header cells. This examination reveals that “53,196,521.18” represents the income from the “main business” project.
@@ 4.2 Stage 1: Self-supervised Instruction Tuning
Illustrated in Fig. 2, HGT utilizes the tabular HG encoder’s output as soft prompts Li and Liang (2021), which form part of the LLM input. The weights of both modules are tuned through self-supervised instruction tuning to align the vector representation spaces of the two modules. This subsection provides a detailed description of the training process, which involves three granularity self-supervised tasks.
4.2.1 Tabular HG Encoder
Following the conversion of the table into a tabular HG, HGT introduces an RGNN Wang et al. (2020a) as the encoder for tabular HG data. The encoder takes a tabular HG as input and generates vector representations for the tabular nodes as output. The RGNN employs a message-passing mechanism to collect semantic and topological information from neighboring cell nodes, considering various edge types. Typically, the RGNN comprises multiple graph aggregation layers, with each layer using the following aggregation formula for each cell node $v$:
$h_{v}^{(l+1)}={\sigma}\left(\sum_{e_{t}\in E_{t}}\sum_{u\in N_{v}^{e_{t}}}\frac{1}{\left|N_{v}^{e_{t}}\right|}W_{e_{t}}^{(l)}h_{u}^{(l)}\right),$ (1)
where $l$ is the $l^{th}$ layer, $N_{v}^{e_{r}}$ means all neighbor cell nodes of $v$, $E_{r}$ is the set of edge type, $\sigma$ is an activation function and $W_{e_{t}}^{(l)}$ is trainable weights. The last hidden representation is used as part of the input for subsequent modules in the framework.
4.2.2 Multi-granularity Self-supervised Instruction Tasks
Drawing inspiration from the methodologies presented in LLaVA Liu et al. (2023), Time-LLMJin et al. (2023a), and GraphGPT Tang et al. (2023), we develop a method to effectively align the vector spaces of two distinct modalities of data: tables and NLs. We introduce the soft prompts technique as a bridge between the two encoders and focus on fine-tuning only the tabular graph encoder and the LoRAHu et al. (2021) module within the LLM. This process, designed to be lightweight, enables the LLM to grasp the topological information of tables through semantic instruction learning.
@@ 4.2.2 Multi-granularity Self-supervised Instruction Tasks
Instruction learning Wang et al. (2022); Ouyang et al. (2022a), a technique that merges fine-tuning with prompt learning, significantly enhances the generalization capabilities of LLMs. Therefore, before the model-tuning process, we pre-format the self-supervised training data into an instruction-based QA format. Examples of training data for the three tasks—Table Row Classification, Table Cell Matching, and Table Context Generation—are illustrated in Fig. 4.
@@
To align the LLM with the HG encoder, the HGT incorporates additional tokens into the vocabulary: <tabular_node>, <table_start>, and <table_end>. The token <tabular_node> serves as a placeholder for table tasks within the instruction text, allowing for the substitution with actual table node vectors post-processing by the LLM’s Embedding layer. The quantity of these placeholders matches the number of table nodes relevant to the current task. For instance, since there are 10 cell tokens in Fig. 4, the <tabular_node> is replicated 10 times. The tokens <table_start> and <table_end> signify the beginning and ending delimiters of the table placeholders.
@@
The forward propagation in HGT begins with the input of a <Table, Instruction> pair, where the instruction text passes through the LLM’s Embedding layer, and the table is transformed into a Tabular HG before being processed by the RGNN. The LLM Embedding layer assigns embeddings to each token, creating a sequence, while the RGNN’s output provides aggregated vector representations for each table node. Subsequently, HGT replaces the embeddings corresponding to the table placeholders within the sequence with the actual node vectors. This adjusted embedding sequence is then fed into the LLM’s remaining layers.
@@
To enhance the HGT understanding of tables, we innovatively devise three tasks tailored for the graph-enhanced LLM, each varying in granularity: Table Row Classification, Table Cell Matching, and Table Context Generation. The complexity of these tasks increases progressively, aiming to elevate the model’s capability to comprehend semantic information across different levels. This tiered method is designed to systematically broaden the model’s semantic awareness. Below is the detailed descriptions of each task.
@@
Table Row Classification (TRC). The objective of this task is to train the model to accurately identify the category of each table row. Utilizing the vector representations of Row nodes provided by the HG encoder, the model discerns whether a node corresponds to a Header Row or a Data Row. The categorization of Row nodes is automated by an algorithm elaborated in Appendix. B, eliminating the need for manual intervention and thus qualifying as a self-supervised training process. In the dataset we utilize, the labeling accuracy of the algorithm reaches 97.6%. This task facilitates the model’s initial grasp of coarse-grained information about the structure of table rows.
@@
Table Cell Matching (TCM). The task involves supplying the model with vector representations of each cell node alongside a list of shuffled cell contents. The model’s objective is to correctly pair each cell node vector with its corresponding original text. For instance, referring to the table depicted in Fig. 3, the model needs to align the node vector of $c_{1,1}$ with the string "income" within the list correctly. This training task enables the model to discern the semantic information encapsulated within the cell contents based on the graph node vectors. Essentially, it aligns the semantic space of the HG encoder and the LLM.
@@
Table Context Generation (TCG). The objective of this task is to enable the model to generate context surrounding table data, utilizing the vector representations from the Table node. This task facilitates the model’s learning of the table’s global semantic information, proving beneficial for tasks requiring a comprehensive understanding of the table as an entity.
Examples of data sets in instruction format for the three tasks are shown in Fig. 4.
4.3 Task-specific Instruction Tuning

### v2

# HeGTa: Leveraging Heterogeneous Graph-enhanced Large Language Models for Few-shot Complex Table Understanding

@@ Abstract
Table Understanding (TU) has achieved promising advancements, but it faces the challenges of the scarcity of manually labeled tables and the presence of complex table structures. To address these challenges, we propose HeGTa, a heterogeneous graph (HG)-enhanced large language model (LLM) designed for few-shot TU tasks. This framework aligns structural table semantics with the LLM’s parametric knowledge through soft prompts and instruction tuning. It also addresses complex tables with a multi-task pre-training scheme by incorporating novel multi-granularity self-supervised HG pre-text tasks. We empirically demonstrate the effectiveness of HeGTa, showing that it outperforms the SOTA for few-shot complex TU on several benchmarks.
@@ 1 Introduction
Table Understanding (TU) seeks to learn informative embeddings of tables containing inherently tabular semantics, often formatted in ways not easily interpretable by machines, as shown in Fig. 1. This endeavor enhances machine performance across a range of table-related tasks, such as Table QA (Wang et al. 2024), Cell Type Classification (Jia et al. 2023) and Table Type Classification (Lu et al. 2024).
@@
Yet, in real-world scenarios, TU faces the challenges of a lack of sufficient human annotations and the presence of complex table structures, which diminishes the effectiveness and applicability of existing frameworks. As for the first challenge, the data-hungry nature of existing frameworks results in diminished performance in few-shot TU scenarios, where only several samples are annotated. Although some studies (Dong et al. 2022; Herzig et al. 2020) have adopted pre-training with encoder-only architectures (Devlin et al. 2019) to alleviate the annotation burden, these solutions still require considerable amounts of labeled data for task-specific fine-tuning. Regarding the second challenge, while current frameworks attempt to capture structural information in tables through the use of position embeddings (Wang et al. 2020c) or by modeling tables as graphs (Jia et al. 2023; Wang et al. 2021), these methods are effective for simple tables, but wane with complex tables. This shortfall arises because the cell-to-cell relationships in complex tables are more intricate than those in simple tables, as shown in Fig. 1.
@@
Figure 1: Few-shot complex Table Understanding. Complex tables contain intricate cell-to-cell relationships, including dependency, hierarchical, and parallel ones.
@@
Fortunately, recent advancements have introduced techniques that can be integrated into the TU framework to tackle the challenges: 1) Large Language Models (LLMs) have shown remarkable effectiveness in managing few-shot tasks involving data in other modalities, such as vision (Liu et al. 2023) and time series (Jin et al. 2023a). This is achieved through instruction tuning and soft prompts, which seek to align the semantic spaces of the other modality encoder with that of the LLM, a method that can also be applied to tabular data. 2) Self-supervised heterogeneous graph (HG) pre-training (Ma, Liu, and Zuo 2023; Yang et al. 2022) enables models to navigate multifaceted relationships within data, such as complex intercell relationships in Fig. 1, using substantial volumes of unlabeled data.
@@
We propose HeGTa, an HG-enhanced LLM framework for the few-shot complex TU. To enable HeGTa to model intricate relationships within complex tables, HeGTa begins with modeling the tabular data with the HG and then processing it through a tabular graph encoder to generate vectors imbued with structural information. Additionally, to ensure that the LLM achieves comparable performance in table tasks to its performance in natural language (NL) tasks, especially in few-shot scenarios, we align the representation spaces of the tabular graph encoder and the LLM via instruction tuning. Specifically, we integrate table vectors containing structural information as soft prompts within the LLM inputs and innovatively design three multi-granularity self-supervised tasks tailored for both tables and LLMs for pre-training. After pre-training specific parameters of HeGTa via these self-supervised tasks, the HeGTa can capitalize on the LLM’s exceptional generalization capability to adapt to downstream tasks with minimal data samples. To validate the performance of HeGTa in few-shot complex TU, we conduct extensive comparative experiments with existing powerful baselines on nine publicly available datasets for three specific table-related tasks in few-shot scenarios. The experimental results show that HeGTa exceeds the current SOTA for the few-shot complex TU across multiple benchmark datasets. In summary, our main contributions are as follows.
@@
  • To the best of our knowledge, we are the first to propose a framework to align table semantics with NLs’ to empower the LLM to perform on tables with the same few-shot proficiency as it does with NLs.
@@
  • We conduct extensive experiments on nine publicly available datasets, and the experimental results show that HeGTa exceeds the current SOTA for few-shot complex TU across multiple benchmark datasets.
@@ 2.1 Graph-based Table Understanding
Many studies (Du et al. 2021; Jin et al. 2023b) have converted tables into graphs and utilized graph encoders to capture tables’ inherent topological information. These frameworks include employing homogeneous graphs or utilizing a basic node-linking strategy that connects cells exclusively to their adjacent counterparts. Consequently, such frameworks underperform when dealing with complex table structures.
2.2 LLM-based Textual Table Understanding
Following the success of LLMs in NL tasks, some efforts (Zhang et al. 2023a; Chen 2022; Sui et al. 2024) have extended their application to table-related tasks. Although these frameworks leverage the exceptional generalization capacity of LLMs to achieve SOTA performance in some few-shot tasks, they resort to simply converting the table into a row-by-row NL format as the input. This process leads to a loss of the tables’ intrinsic topological information.
2.3 LLM-based Symbolic Table Understanding
Inspired by Program-of-Thoughts(Chen et al. 2022), several studies (Cao et al. 2023; Cheng et al. 2022; Wang et al. 2024; Zhang et al. 2023b; Ye et al. 2023) have adapted semantic parsing techniques, traditionally applied to relational database tables, to general Table QA tasks. This involves transforming the table into a format interpretable by query languages, such as SQL, and subsequently utilizing LLMs to generate a symbolic program to retrieve the answer. This methodology represents the current SOTA in Table QA. However, its architectural limitations restrict its applicability to other table-related tasks.
@@ 2.4 Bert-like Encoder-only Table Understanding
Since the rise in popularity of pre-training models like BERT (Devlin et al. 2019), there has been considerable effort devoted to designing specialized encoding methods for tabular data and unique pre-text objectives for pre-training (Yin et al. 2020; Cheng et al. 2021a; Jia et al. 2023). Despite their utilization of self-supervised training, these methods still require a substantial amount of labeled data during fine-tuning for downstream tasks. Additionally, while they incorporate positional embedding into serialized tabular data, they do not effectively capture topological information.
@@ 3 Task Definition
Given a table $T=\{c_{i,j}|0\leq i<N,0\leq j<M\}$ where $N$ is the number of rows, $M$ is the number of columns, and $c_{i,j}$ is the cell located in the $i^{th}$ row and $j^{th}$ column. Merged cells, characterized by a row span or column span greater than 1, are prevalent in complex tables. For instance, the cell labeled ”project” in the top-left corner of the table shown in Fig. 3 has a row span of 2. We assign the coordinates of such merged cells based on the position of the top-left cell before merging. Hence, the coordinate of the ”project” cell is designated as (0,0).
Specific sub-TU tasks for Evaluation
@@
Table QA (TQA) demands that the model produce an answer $y_{a}$ in response to a natural language question $q$, with table $T$ serving as the reference for deriving the answer $y_{a}$.
@@ 4 METHODOLOGY
Figure 2: An overview of HeGTa framework. HeGTa processes <table, instruction> as an input. First, the table is converted into an HG and processed by a Tabular HG encoder to generate a vector for each tabular node, while the LLM transforms instruction texts into initial token embeddings. Subsequently, the HG encoder’s outputs serve as soft prompts for the LLM, enabling the replacement of placeholder embeddings with actual tabular node vectors. The modified embedding sequence is then processed by the remaining LLM layers. Throughout Stage 1 and Stage 2, only the weights of red components are tuned.
This section explains three phases of HeGTa: tabular HG construction and two tuning stages, as shown in Fig. 2.
@@ 4.1 Tabular Heterogeneous Graph Construction
Figure 3: Table-to-heterogeneous graph conversion. Node types are color-coded: Table (green), Row (red), Data Cell (blue), and Header Cell (yellow). Edge types are similarly color-coded, with bidirectional edges shown as undirected lines. Some edges are omitted for clarity.
@@
Given that HGs are more proficient at capturing diverse relationships compared to homogeneous graphs, we employ HGs with heuristic node-linking rules to model the structure of complex tables effectively. The subsequent subsections detail the process of converting tabular data into HGs, including the creation of nodes and the heuristic rules for establishing edges between nodes. The process of creating a tabular HG is shown in Fig. 3.
Four Tabular Node Types
@@
TABLE, ROW, Header CELL, and Data CELL nodes are denoted as green, red, yellow, and blue nodes in Fig. 3. 1) The TABLE node represents the content described by the table, facilitating the table-level tasks. 2) The ROW node signifies the information contained within a row, aiding in the prediction of the row’s type during self-supervised training. 3) The Header CELL node denotes cells located in the header row, identifying column schemas or categories. 4) The Data CELL node represents cells in data rows, meaning the actual data entries of the table.
Creating Tabular Nodes
@@
First, CELL nodes are created for each cell in the table, with each node denoted as $c_{i,j}$, where $i$ and $j$ represent the cell’s coordinates in the original table. If a CELL node is located in the table’s header section or is part of a merged cell spanning the entire table width, it is viewed as a Header Cell; otherwise, it is identified as a Data Cell. The algorithm for determining the header section is elaborated in Appendix. Subsequently, ROW nodes are created to match the number of rows in the table, along with a single TABLE node.
Initializing Tabular Node Embeddings
@@
An initialization vector is required for each node in the HG. For Header and Data CELL nodes, we employ the output of S-BERT (Reimers and Gurevych 2019) applied to the text within each cell to obtain their initialization vectors. In the case of ROW nodes $r_{i}$, we initialize a vector by concatenating the text from the cells in the $i^{th}$ row and inputting this concatenated text into the S-BERT. TABLE nodes serve to represent the table’s content itself. In a human’s view, a table’s content can be inferred by examining the cells in the headers. Consequently, we opt to concatenate the text from the cells in the headers and to obtain the embedding.
Adding Edges
@@
To enhance the machine’s understanding of the table’s semantics, we develop the following heuristic rules to link nodes: 1) A TABLE node, representing the whole table, should be linked to all CELL nodes to encapsulate the global semantics. 2) A ROW node derives information from cells within the same row, so the ROW node $r_{i}$ should be linked to each CELL node $c_{i*}$. 3) A strong correlation exists between the semantics of Data CELLs and their corresponding Header CELLs within the same column, so they should be linked. 4) Data CELLs located in the same column exhibit a stronger relational bond compared to those in different columns. Consequently, adjacent Data CELLs within a column are interconnected. 5) All Header CELLs should be interconnected because the interpretation of relationships between Data CELLs across columns necessitates the semantic understanding of their respective Header CELLs. For instance, as demonstrated in Fig. 3, determining the relationship between cell $c_{2,0}$ and cell $c_{2,1}$ requires an examination of their Header CELLs. This examination reveals that “53,196,521.18” represents the income from the “main business” project.
@@ 4.2 Stage 1: Self-supervised Instruction Tuning
Illustrated in Fig. 2, HeGTa utilizes the tabular output of the HG encoder as soft prompts (Li and Liang 2021), which are part of the LLM input. The weights of both modules are tuned through self-supervised instruction tuning to align the vector representation spaces of the two modules. This subsection provides details of the pre-training process, which involves three different granularity self-supervised tasks.
Tabular HG Encoder
Following the conversion of the table into a tabular HG, HeGTa introduces a heterogeneous GNN (HGNN) (Hu et al. 2020) as the encoder for tabular HG data. The encoder takes a tabular HG as input and generates vector representations for the tabular nodes as output. The HGNN employs a message-passing mechanism to collect semantic and topological information from neighboring cell nodes within each layer, considering various edge types.
Multi-granularity Self-supervised Instruction Tasks
Drawing inspiration from the methodologies in LLaVA (Liu et al. 2023), Time-LLM (Jin et al. 2023a), and GraphGPT (Tang et al. 2023), we develop a method to effectively align the vector spaces of two distinct modalities of data: tables and NLs. We introduce the soft prompt technique as a bridge between the two encoders and focus on fine-tuning the tabular graph encoder and the LoRA (Hu et al. 2021) module within the LLM. This process, designed to be lightweight, enables the LLM to grasp the topological information of tables through semantic instruction tuning.
@@
Instruction tuning (Wang et al. 2022; Ouyang et al. 2022), a technique that merges fine-tuning with prompt learning, significantly enhances the generalization capabilities of LLMs. Therefore, before the model-tuning process, we pre-format the self-supervised training data into an instruction-based QA format. Examples of training data for instruction tuning are illustrated in Fig. 4.
@@
To align the LLM with the HG encoder, the HeGTa incorporates additional tokens into the vocabulary: <tabular_node>, <table_start> and <table_end>. The token <tabular_node> serves as a placeholder for table tasks within the instruction text, allowing for the substitution with actual table node vectors post-processing by the LLM’s Embedding layer. The quantity of these placeholders matches the number of table nodes relevant to the current task. The tokens <table_start> and <table_end> signify the beginning and ending delimiters of the table placeholders.
@@
The forward propagation in HeGTa begins with the input of a <Table, Instruction> pair, where the instruction text passes through the LLM’s embedding layer, and the table is transformed into a Tabular HG before being processed by the HGNN. The LLM embedding layer assigns embeddings to each token, creating a sequence, while the HGNN’s output provides aggregated vector representations for each table node. Subsequently, HeGTa replaces the embeddings corresponding to the table token placeholders within the sequence with the actual node vectors. This adjusted embedding sequence is then fed into the LLM’s remaining layers.
@@
To enhance the HeGTa’s understanding of tables, we introduce three tasks tailored for the graph-enhanced LLM, each varying in granularity: Table Row Classification, Table Cell Matching, and Table Context Generation. These tasks are designed to integrate table representations into the LLM’s semantic space. The complexity of these tasks increases progressively, aiming to incrementally improve the model’s ability to comprehend semantic information. Detailed descriptions of each task are provided below.
@@
Table Row Classification (TRC). The objective of this task is to train the model to accurately identify the category of each table row. Utilizing the vector representations of Row nodes provided by the HG encoder, the model discerns whether a node corresponds to a Header Row or a Data Row. The categorization of Row nodes is automated by an algorithm elaborated in Appendix, eliminating the need for manual intervention and thus qualifying as a self-supervised training process. In the dataset we utilize, the labeling accuracy of the algorithm reaches 97.6%. This task facilitates the model’s initial grasp of coarse-grained information about the structure of table rows.
@@
Table Cell Matching (TCM). The task involves supplying the model with vector representations of each cell node alongside a list of shuffled cell contents. The model’s objective is to correctly pair each cell node vector with its corresponding original text. For instance, referring to the table depicted in Fig. 3, the model needs to align the node vector of $c_{1,1}$ with the string ”income” within the list correctly. This training task enables the model to discern the semantic information encapsulated within the cell contents based on the graph node vectors. Essentially, it aligns the semantic space of the HG encoder and the LLM.
@@
Table Context Generation (TCG). This task aims to enable the model to generate context surrounding table data, utilizing the vector representations from the Table node. This task facilitates the model’s learning of the table’s global semantic information, proving beneficial for tasks requiring a comprehensive understanding of the table as an entity.
Examples of datasets in instruction format for the three tasks are provided in Fig. 4.
4.3 Stage 2: Task-specific Instruction Fine-tuning